OpenStudy (anonymous):

i need help figuring out how to work this problem. the width of a rectangle is 6 in less than the lengh. the perimeter is 120 in. find the dimensions

OpenStudy (anonymous):

let the length be x

OpenStudy (anonymous):

then the width is x-6 the perimeter is 2(leng + breadth) = 120 2 (x + x-6) = 120 2x-6 = 60 x=33

OpenStudy (anonymous):

ok ty so much

OpenStudy (anonymous):

2x+2y = 12 or x+y=60 x=y-6 y + y-6=60 2y=66 y=33 x=27